OpenStudy (anonymous):

The face of a tent is shaped like a triangle with an area of 12 feet squared . The area of the triangle is represented by the formula A equals one half B times H. The height of the tent is 4 feet. How wide is the base of the tent?

OpenStudy (anonymous):

Rules to solve word problems 1. Translate words into mathematical terms/symbols and use them to label a geometric drawing 2. Write the resulting equation(s)

OpenStudy (arbershabani97):

A=1/2*b*h A=12, b=?, h=4ft 12=1/2*b*4 12=1/2*4*b 12=2*b b=... if you don't understand anything here ask freely